OpenStudy (anonymous):

Which is an equation in slope-intercept form for the line passing through (–1, 5) and (5, 8).

OpenStudy (anonymous):

It's y= 1/2x + 11/2 right?

OpenStudy (the_fizicx99):

I suck the the quation tab, y2-y1/x2-x1 8-5/5-(-1) 3/6 = .5 y-y1=m(x-x1) y-5 =.5(x+1) y-5=.5x +.5 +5 +5 y = .5x + 5.5

OpenStudy (the_fizicx99):

@its_deadpool_